# POMATO: Marrying Pointmap Matching with Temporal Motions for Dynamic 3D Reconstruction

### v1

@@ Abstract
3D reconstruction in dynamic scenes primarily relies on the combination of geometry estimation and matching modules where the latter task is pivotal for distinguishing dynamic regions which can help to mitigate the interference introduced by camera and object motion. Furthermore, the matching module explicitly models object motion, enabling the tracking of specific targets and advancing motion understanding in complex scenarios. Recently, the proposed representation of pointmap in DUSt3R suggests a potential solution to unify both geometry estimation and matching in 3D space, but it still struggles with ambiguous matching in dynamic regions, which may hamper further improvement. In this work, we present POMATO, a unified framework for dynamic 3D reconstruction by marrying POintmap MAtching with Temporal mOtion. Specifically, our method first learns an explicit matching relationship by mapping RGB pixels from both dynamic and static regions across different views to 3D pointmaps within a unified coordinate system. Furthermore, we introduce a temporal motion module for dynamic motions that ensures scale consistency across different frames and enhances performance in tasks requiring both precise geometry and reliable matching, most notably 3D point tracking. We show the effectiveness of the proposed pointmap matching and temporal fusion paradigm by demonstrating the remarkable performance across multiple downstream tasks, including video depth estimation, 3D point tracking, and pose estimation. Code and models are publicly available at https://github.com/wyddmw/POMATO.
@@ 1 Introduction
Image-based 3D reconstruction is a fundamental task in computer vision with a wide range of applications including SLAM [37], robotics [17, 47], autonomous driving [51], and novel view synthesis [5]. While substantial progress has been achieved in static 3D reconstruction [14, 49, 21, 42, 24], dynamic scenes remain significantly more challenging. The presence of moving objects introduces random motion and deformation, which can interfere with the learning of local structure and camera motion, thus complicating accurate geometry estimation. These scenarios require joint modeling of both scene geometry and object motion. Moreover, downstream tasks such as 3D point tracking demand precise geometry estimation and robust matching across views. To effectively distinguish dynamic regions, it is essential to establish reliable correspondences between frames. Some pioneering works have attempted to address dynamic motion by incorporating auxiliary matching modules, such as optical flow [40, 50] or 2D tracking [45]. However, these approaches may suffer from domain gaps and accumulated errors between modules, limiting their effectiveness. A unified framework that seamlessly integrates geometry estimation and matching for dynamic 3D reconstruction remains a critical and underexplored challenge.
@@
Recently, DUSt3R [42] proposes a promising solution to address this challenge. It introduces the concept of a pointmap that assigns each pixel in an image to a corresponding 3D coordinate. The network utilizes a standard transformer-based encoder-decoder architecture and receives a pair of images as input. The system incorporates two parallel decoders with a regression head to predict pointmaps for each view within the same coordinate system as the initial view. However, the representation of the pointmap still faces limitations when handling ambiguous correspondence in dynamic scenes, as illustrated in Fig. 2. In static regions, the pointmap of the second image accurately reflects the rigid camera transformation from the first view, where identical RGB pixels correspond to the same 3D coordinates (orange point). Conversely, in dynamic regions, pixel displacements or deformations due to object motion introduce inconsistencies in the predicted pointmaps across views (red point), resulting in inaccurate 3D correspondences. This inconsistency fails to depict the accurate matching relationships in 3D space.
To tackle this problem, we present POMATO, a unified network for dynamic 3D reconstruction by marrying POintmap MAtching with Temporal mOtion. We introduce an auxiliary matching head that leverages decoder tokens from the second image while preserving matching features through iterative cross-attention across views. This module is tasked with predicting the pointmap for the second image, explicitly conditioned on the features of the first view. Specifically, for each pixel in the second image, the predicted pointmap value corresponds to the 3D coordinate of its matching pixel in the first image.
The proposed pointmap matching representation facilitates the establishment of explicit 3D point correspondences, which can be directly utilized for motion analysis. Moreover, we further extend our POMATO to 4D video sequence input by introducing a temporal motion module to enhance the learning of temporal motions which helps to guarantee the scale consistency across different frames and improves performance in tasks where both accurate geometry and reliable matching are paramount, most notably 3D point tracking.
@@
The proposed pointmap-based matching representation enables explicit 3D point correspondences, which can be directly utilized for motion analysis. To further extend our approach, we introduce POMATO for 4D video sequence input by incorporating a temporal motion module. This module enhances the learning of temporal motions with scale consistency across frames, and improves performance on tasks where both accurate geometry and robust matching are essential—most notably, 3D point tracking. Compared with the recent temporal 3D reconstruction methods [41, 39] based on an autoregression manner where the previous frames are blocked from the lately added frames, our temporal motion module is based on the self-attention mechanism along the temporal dimension, facilitating a thorough interaction across all frames. Our POMATO is trained in a two-stage manner. In the first stage, we use pairwise input images to learn fundamental geometry and matching capacity. Then we extend the input to sequential video input and insert the temporal motion module, enabling the model to effectively learn motions along the temporal dimension.
@@
Our contributions can be summarized as three folds: First, we propose a novel approach that unifies the fundamental geometry estimation and motion understanding for dynamic 3D reconstruction into a single network by incorporating the representation of pointmap matching. Second, we introduce a temporal motion module to facilitate the interactions of motion features along the temporal dimension which significantly improves the performance in tasks where both accurate geometry and precise matching are required for video sequential input—most notably, 3D point tracking. Third, we demonstrate promising performance on a range of 3D vision tasks, including video depth estimation on dynamic scenes, 3D point tracking, and camera pose estimation.
@@ 2 Related Work
Geometry estimation refers to the process of determining the spatial properties and structures from different forms of visual data. Direct recovery of 3D geometry from a single RGB image is by nature an ill-posed problem. Many recent works [21, 49, 3, 14] have tried to leverage strong pre-trained models to learn generalizable depthmaps from large-scale real and synthetic datasets to solve ambiguities. For example, Marigold [21], Geowizard [10], and GenPercept [46] aim at leveraging the generative priors from pre-trained diffusion models by finetuning them on synthetic datasets. Depthanything V2 [49] proposes to estimate scale-and-shift invariant disparity map by finetuning DINOV2 [27] model on synthetic datasets and large-scale pseudo labels. Depth Pro [3] further propose a FOV head to estimate the metric depthmap from a single image without relying on camera intrinsics as input. Due to the scale ambiguity in the monocular depth estimation models, ChronoDepth [34], DepthCrafter [15], and Depth-any-video [48] proposes to learn temporal consistent depthmaps by leveraging the priors from a video generative model, i.e. SVD [2]. In another line of the research, multi-view stereo reconstruction (MVS) methods seek to reconstruct visible surfaces from multiple viewpoints. Traditional MVS [11] and SfM pipelines break the reconstruction pipeline into several sub-problems, e.g., feature extraction [7], image matching [1, 25], triangulation, and bundle adjustment [6]. The chain is complicated and accumulates noise for every single step, thus often resulting in unsatisfactory performance in complex real-world scenes. Recognizing the limitations of previous MVS methods, seminal work DUSt3R [42] proposes 3D pointmaps representation, and trains a network from large-scale data to regress the dense and accurate pointmaps from a pair of images. The camera intrinsics and relative camera poses can be implicitly inferred from the two-view pointmaps. However, it still can not handle reconstruction for dynamic scenes.
@@
Motion representation. Optical flow is a commonly used representation for 2D motion, which is defined as a 2D vector field describing the apparent movements of each pixel between a pair of images. RAFT [36] is a representative work for pairwise optical flow estimation, which employs a 4D cost volume and recurrently estimates the optical flow. Some follow-up methods further extend it to multi-frame (3-5 frames) settings, which is still insufficient for long-range tracking. To resolve the problem, Particle Video [33] represent video motion by using a set of particles. Each particle is an image point sample with a long-duration trajectory and other properties. Particle videos have two key advantages over optical flow: (1) persistence through occlusions, and (2) multi-frame temporal context. Some recent works, PIPs [13], TAPIR [8] and Cotracker [20] have renewed interest in this representation and show promising long-term 2D point tracking results. Recognizing the advantage of point representation, SpatialTracker [45] lifts the 2D points into 3D and performs tracking in the 3D space. Though it can handle occlusions and enhance 3D tracking accuracy, it still relies on a separate monocular depth estimator, which prevents it performing 3D point tracking in an end-to-end fashion.
@@
Multi-view dynamic reconstruction. Our work is closely connected to multi-view dynamic 3D reconstruction techniques. Early works [32, 30] take the straightforward idea that first pre-segment the scene into different regions, each corresponding to a single rigid part of an object, then apply the rigid-SfM technique to each of the regions. However, in a general dynamic setting, the task of densely segmenting rigidly moving objects or parts is not trivial. Some of the recent Neural Radiance Fields (NeRF)  [26] and Gaussian Splatting  [22] based methods have achieved state-of-the-art results. However, most of these methods require simultaneous multi-view video inputs or require predefined templates [16]. Shape of motion [40], proposes a new dynamic scene representation to represent the dynamic scene as a set of persistent 3D Gaussians, and optimize the representation from a monocular video by leveraging monocular depth estimation priors and 2D track estimates across frames. MonST3R directly finetuned the original DUSt3R model upon some synthetic datasets that contain dynamic scenes. Our POMATO also represents the scene as the 3D pointmap, Different from MonST3R [50], we propose a point matching head to unify geometry and matching estimation for dynamic reconstruction.
@@ 3 Method
Figure 4: Inference pipelines for point tracking, video depth, and multi-view reconstruction. $t_{k}$ indicates the keyframe in the sequence.
Figure 5: Architecture of our temporal motion module. We insert a transformer-based motion module into the vanilla DPT [31] head at different feature scales to enhance the consistency along the temporal dimension.
@@ 3.1 Preliminary
The overview of our POMATO is demonstrated in Fig.3. We inherit the definition of pointmap ${\bf X}\in{\mathbb{R}}^{H\times W\times 3}$ in DUSt3R [42] as a dense 2D field of 3D points which maps to its corresponding RGB pixels. Given a pair of input images ${\bf I}^{1},{\bf I}^{2}\in{\mathbb{R}}^{H\times W\times 3}$ from two different views, a weight-sharing ViT first extracts the corresponding features ${\bf F}^{1},{\bf F}^{2}$ for each view. Two parallel branches are then introduced to decode the geometric structures and enhance the feature alignment via cross-attention in decoder modules, following a regression head to estimate pointmaps ${\bf X}^{1,1},{\bf X}^{2,1}\in{\mathbb{R}}^{H\times W\times 3}$ along with a confidence map ${\bf C}^{1,1},{\bf C}^{2,1}\in{\mathbb{R}}^{H\times W}$ for each image view. Generally, ${\bf X}^{n,m}$ indicates the pointmap ${\bf X}^{n}$ from camera $n$ expressed in camera $m$’s coordinate frame, which is obtained by a rigid transformation:
@@
The task for Decoder 1 and its regression head can be briefly summarized as estimating the 3D points for ${\bf I}^{1}$ in its own coordinate system while Decoder 2 and its regression head are responsible for estimating pixel-wise 3D coordinates for ${\bf I}^{2}$ in ${\bf I}^{1}$’s coordinate system after a rigid transformation of global rotation and translation. In the following contents, we will first introduce our POMATO with pairwise input images and then extend it to the video sequence input with our temporal motion module.
@@ 3.2 Pointmap Matching with Pairwise Input
$\mathcal{L}_{\text{m}}=\left\|\frac{1}{z}{{\bf X}^{2,1}_{m}}-\frac{1}{\bar{z}}{\bar{{\bf X}}^{2,1}}_{m}\right\|,\vspace{-1mm}$ (3)
where $\bar{{\bf X}}^{2,1}_{m}$ is the ground truth pointmap matching which can be obtained following Eq. 2 on 2D tracking dataset along with the depth and camera information. $z,\bar{z}$ are the same norm factor defined in DUSt3R. The matching confidence ${\bf C}^{2,1}_{m}$ is also jointly learned with the confidence loss for Head1 and Head2 within valid regions.
$\mathcal{L}_{\text{mconf}}={\bf C}^{2,1}_{m}\mathcal{L}_{\text{m}}-\alpha\text{log}{\bf C}^{2,1}_{m}\vspace{-1mm}$ (4)
@@ 3.3 Dynamic Mask Estimation
Taking advantage of the explicit pointmap matching head, our POMATO can directly perform dynamic mask estimation without introducing a third assistant module like the optical flow model, getting rid of the additional computation cost and the potential domain gap. For an image pair $\{{\bf I}^{i},{\bf I}^{j}\}$ along with the estimation of ${\bf X}^{j,i}$ from Head2 and ${\bf X}^{j,i}_{m}$ from Head3, the dynamic mask ${\bf D}^{j,i}$ can be obtained by comparing the difference between ${\bf X}^{j,i}$ and ${\bf X}^{j,i}_{m}$:
${\bf D}^{j,i}=||{\bf X}^{j,i}_{m}-{\bf X}^{j,i}||>\alpha,$ (5)
where $\alpha$ is a dynamic threshold defined as $3\times\text{median}(||{\bf X}^{j,i}_{m}-{\bf X}^{j,i}||)$. The explicit dynamic mask can be incorporated into the global alignment process to minimize the interference of moving objects for pose estimation and 3D reconstruction. The incorporation of dynamic masks for improving global alignment is detailed in the supplementary materials.
3.4 Motion Module for Video Pointmap Matching
@@
With the fundamental capacity of geometric estimation and pointmap matching for pairwise images, we extend our POMATO to 4D video sequences for temporal-related tasks by inserting a transformer-based motion module into the vanilla DPT [31] head and we refer this enhanced regression head as the ”temporal DPT head”. The architecture of the temporal DPT head is illustrated in Fig. 5. For a set of decoder tokens ${\bf G}\in{\mathbb{R}}^{B,T,N,C}$ where $B,T,N,C$ represent the batch size, window length of a video sequence, token number, and token dimension, respectively, we merge the token number dimension into the batch axis and apply the motion module which consists of two blocks of standard multi-head self-attention modules and feed-forward networks along the temporal dimension $T$. As shown in Fig. 5, the motion module is only applied to decoder tokens of low feature resolutions in terms of minimizing the GPU memory consumption. In this work, the window length $T$ for a video sequence is set to 12 as default. Explorations of different window lengths can be found in Sec. 4. When training with the motion module, we freeze parameters in both the encoder and decoder and finetune the temporal DPT head only. With the introduction of the temporal DPT head, our POMATO can be introduced to applications on the dynamic videos.
3D Point Tracking. Given a video sequence of $T$ frames $\{{\bf I}^{t_{1}},{\bf I}^{t_{2}},...,{\bf I}^{t_{T}}\}$, we can create a set of stereo image pairs: $\{({\bf I}^{t_{k}},{\bf I}^{t_{2}}),({\bf I}^{t_{k}},{\bf I}^{t_{3}}),...,({\bf I}^{t_{k}},{\bf I}^{t_{T}})\}$, where ${\bf I}^{t_{k}}$ is the keyframe. For each pair $({\bf I}^{t_{k}},{\bf I}^{t_{n}})$, where $n\in\{1,2,...T\}$, we compute the dense pointmap matching ${\bf X}^{t_{k},t_{n}}_{m}$ for each reference pixel in the coordinate of the keyframe. As shown in the top part of Fig.4, all the target frames are fed into the Head1 for estimating the pointmap under their own coordinate system while the grouped keyframes are fed to the Head3 for finding the corresponding points in each target frame. The introduced motion module facilitates the interaction along the temporal dimension, enhancing the scale consistency for the global pointmap matching results. Once all pairwise pointmap matching is completed (which can be parallelized using batch operations), we obtain a set of dense 3D tracking results in terms of the keyframe: $\{{\bf X}^{t_{k},t_{2}}_{m},{\bf X}^{t_{k},t_{3}}_{m},...{\bf X}^{t_{k},t_{T}}_{m}\}$. For sparse 3D tracking, the dense tracking results can be sparsified by indexing with the 2D coordinates of query points in ${\bf I}^{t_{k}}$. When tracking across a video sequence longer than $T$ frames, we use a simple sliding-window approach with an overlap of 4 frames for a window length of 12 frames. For the new appearing frames, we directly compute the tracking to the target keyframe. For the overlapped frames, we use a linear weighted method to fuse the prediction from the last sequence to the current. Besides the default loss function for Head1 and Head3, we also introduce a temporal consistency loss by computing the scale factor along the temporal dimension. For pointmap matching loss in Eq.4, the additional temporal tracking loss $\mathcal{L}_{\text{t}}$ is:
$\mathcal{L}_{\text{t}}=\frac{1}{T}\sum^{T}_{i=1}\left\|\frac{1}{z_{T}}{{\bf X}^{t_{k},t_{i}}_{m}}-\frac{1}{\bar{z}_{T}}{\bar{{\bf X}}^{t_{k},t_{i}}}_{m}\right\|,\vspace{-1mm}$ (6)
where the scaling factors $z_{T}=$ norm $({\bf X}^{t_{k},t_{1}}_{m},...,{\bf X}^{t_{k},t_{T}}_{m})$ and $\bar{z}_{T}=$ norm $(\bar{{\bf X}}^{t_{k},t_{1}}_{m},...,\bar{{\bf X}}^{t_{k},t_{T}}_{m})$. The additional temporal loss is similarly applied to the Head1.
Video Depth Estimation. As shown in the middle part of the Fig. 4, the input video sequence is formulated to a set of identical image pairs $\{({\bf I}^{t_{1}},{\bf I}^{t_{1}}),({\bf I}^{t_{2}},{\bf I}^{t_{2}}),...,({\bf I}^{t_{T}},{\bf I}^{t_{T}})\}$. The video depth estimation task involves both Head1 and Head2 where the predictions from each head are identical. We use the output of Head1 as our final depth estimation. Similarly, the temporal consistency loss is also applied to both heads.

### v2

@@ Abstract
Recent approaches to 3D reconstruction in dynamic scenes primarily rely on the integration of separate geometry estimation and matching modules, where the latter plays a critical role in distinguishing dynamic regions and mitigating the interference caused by moving objects. Furthermore, the matching module explicitly models object motion, enabling the tracking of specific targets and advancing motion understanding in complex scenarios. Recently, the proposed representation of pointmap in DUSt3R suggests a potential solution to unify both geometry estimation and matching in 3D space, effectively reducing computational overhead by eliminating the need for redundant auxiliary modules. However, it still struggles with ambiguous correspondences in dynamic regions, which limits reconstruction performance in such scenarios. In this work, we present POMATO, a unified framework for dynamic 3D reconstruction by marrying POintmap MAtching with Temporal mOtion. Specifically, our method first learns an explicit matching relationship by mapping RGB pixels across different views to 3D pointmaps within a unified coordinate system. Furthermore, we introduce a temporal motion module for dynamic motions that ensures scale consistency across different frames and enhances performance in 3D reconstruction tasks requiring both precise geometry and reliable matching, most notably 3D point tracking. We show the effectiveness of our proposed POMATO by demonstrating the remarkable performance across multiple downstream tasks, including video depth estimation, 3D point tracking, and pose estimation. Code and models are publicly available at https://github.com/wyddmw/POMATO.
@@ 1 Introduction
Image-based 3D reconstruction is a fundamental task in computer vision with a wide range of applications including SLAM [39], robotics [19, 49], autonomous driving [53], and novel view synthesis [5]. While substantial progress has been achieved in static 3D reconstruction [16, 51, 23, 44, 26], dynamic scenes remain a major hurdle due to complexities like non-rigid motion and deformation, which may hamper the learning of local structure and camera motion, thereby complicating accurate 3D reconstruction for dynamic scenes. These scenarios require explicit modeling of both scene geometry and object motion. Moreover, downstream reconstruction tasks, such as 3D point tracking, demand precise geometry estimation and robust matching across views. To effectively distinguish dynamic regions, it is essential to establish reliable correspondences between different frames. Some pioneering works have attempted to address dynamic motion by incorporating additional auxiliary matching modules, such as optical flow [42, 52] or 2D tracking [47]. However, these approaches may suffer from domain gaps and accumulated errors between modules, limiting their effectiveness. A unified framework that seamlessly integrates geometry estimation and matching for dynamic 3D reconstruction remains a critical and underexplored challenge.
@@
Recently, DUSt3R [44] proposes a promising solution to address this challenge. It introduces the concept of a pointmap that assigns each pixel in an image to a corresponding 3D coordinate. The network utilizes a standard transformer-based encoder-decoder architecture and receives a pair of images as input. The system incorporates two parallel decoders to predict pointmaps for each view within the same coordinate system. However, this representation is limited to static matching and struggles in dynamic scenes, as illustrated in Fig. 2.
@@
To address this problem, we present POMATO, a unified network for dynamic 3D reconstruction by marrying POintmap MAtching with Temporal mOtion. We argue that with iterative cross-attention modules across different views, matching features are well preserved in the decoder tokens. We thus introduce an auxiliary pointmap matching head to learn explicit correspondences. Specifically, for each pixel in the second view, the pointmap matching head predicts the corresponding 3D coordinates of its counterpart in the first view, under the shared coordinate system. Our proposed pointmap-based matching representation enables the establishment of explicit correspondences in 3D space, which can be directly leveraged for motion analysis, especially the estimation of dynamic regions. Moreover, we further extend our POMATO to handle 4D video sequences by introducing a temporal motion module that enhances the learning of temporal motions. This motion module promotes scale consistency across different frames and improves performance in tasks where both accurate geometry and reliable matching are paramount, most notably 3D point tracking. Compared with recent temporal 3D reconstruction methods [43, 41] based on an autoregression manner where the previous frames are blocked from the recently added frames, our temporal motion module is based on the self-attention mechanism along the temporal dimension, facilitating a comprehensive interaction across all frames. Our POMATO is trained in a two-stage manner. In the first stage, we used pairwise input images to learn fundamental geometry and matching capabilities. In the second stage, we extend the input to sequential video frames and incorporate the temporal motion module, enabling the model to effectively capture motions over time.
@@
Our contributions can be summarized in threefold: First, we propose a novel approach that unifies the fundamental geometry estimation and motion understanding for dynamic 3D reconstruction into a single network by incorporating the representation of pointmap matching. Second, we introduce a temporal motion module to facilitate the interactions of motion features along the temporal dimension, which significantly improves the performance in tasks where both accurate geometry and precise matching are required for video sequential input, most notably 3D point tracking. Third, we demonstrate promising performance on 3D vision tasks, including video depth estimation, 3D point tracking, and camera pose estimation.
@@ 2 Related Work
Geometry estimation refers to the process of determining the spatial properties and structures from different forms of visual data. Direct recovery of 3D geometry from a single RGB image is by nature an ill-posed problem. Many recent works [23, 51, 3, 16] have tried to leverage strong pre-trained models to learn generalizable depthmaps from large-scale real and synthetic datasets to solve ambiguities. For example, Marigold [23], Geowizard [11], and GenPercept [48] aim at leveraging the generative priors from pre-trained diffusion models by finetuning them on synthetic datasets. Depthanything V2 [51] proposes to estimate scale-and-shift invariant disparity map by finetuning DINOV2 [29] model on synthetic datasets and large-scale pseudo labels. Depth Pro [3] further propose a FOV head to estimate the metric depthmap from a single image without relying on camera intrinsics as input. Due to the scale ambiguity in the monocular depth estimation models, ChronoDepth [36], DepthCrafter [17], and Depth-any-video [50] proposes to learn temporal consistent depthmaps by leveraging the priors from a video generative model, i.e. SVD [2]. In another line of the research, multi-view stereo reconstruction (MVS) methods seek to reconstruct visible surfaces from multiple viewpoints. Traditional MVS [12] and SfM pipelines break the reconstruction pipeline into several sub-problems, e.g., feature extraction [8], image matching [1, 27], triangulation, and bundle adjustment [7]. The chain is complicated and accumulates noise for every single step, thus often resulting in unsatisfactory performance in complex real-world scenes. Recognizing the limitations of previous MVS methods, seminal work DUSt3R [44] proposes 3D pointmaps representation, and trains a network from large-scale data to regress the dense and accurate pointmaps from a pair of images. The camera intrinsics and relative camera poses can be implicitly inferred from the two-view pointmaps. However, it still can not handle reconstruction for dynamic scenes. MonST3R [52] directly finetuned the original DUSt3R model upon synthetic datasets that contain dynamic scenes.
@@
Motion representation. Optical flow is a commonly used representation for 2D motion. RAFT [38] is a representative work for pairwise optical flow estimation, which employs a 4D cost volume and recurrently estimates the optical flow. Some follow-up methods further extend it to multi-frame (3-5 frames) settings, which is still insufficient for long-range tracking. To resolve the problem, Particle Video [35] represent video motion by using a set of particles. Each particle is an image point sample with a long-duration trajectory and other properties. Particle videos have two key advantages over optical flow: (1) persistence through occlusions, and (2) multi-frame temporal context. Some recent works, PIPs [15], TAPIR [9] and Cotracker [22] have renewed interest in this representation and show promising long-term 2D point tracking results. Recognizing the advantage of point representation, SpatialTracker [47] lifts the 2D points into 3D and performs tracking in the 3D space. Though it can handle occlusions and enhance 3D tracking accuracy, it still relies on a separate monocular depth estimator, which prevents it performing 3D point tracking in an end-to-end fashion.
@@
Multi-view dynamic reconstruction. Our work is closely connected to multi-view dynamic 3D reconstruction techniques. Early works [34, 32] take the straightforward idea that first pre-segment the scene into different regions, each corresponding to a single rigid part of an object, then apply the rigid-SfM technique to each of the regions. Some of the recent Neural Radiance Fields (NeRF)  [28] and Gaussian Splatting  [24] based methods have achieved state-of-the-art results. However, most of these methods require simultaneous multi-view video inputs or require predefined templates [18]. Shape of motion [42], proposes a new dynamic scene representation to represent the dynamic scene as a set of persistent 3D Gaussians, and optimize the representation from a monocular video by leveraging monocular depth estimation priors and 2D track estimates across frames.
@@ 3.1 Preliminary
The overview of our POMATO is demonstrated in Fig.3. We adopt the definition of pointmap ${\bf X}\in{\mathbb{R}}^{H\times W\times 3}$ in DUSt3R [44] as a dense 2D field of 3D points where each point corresponds to its respective RGB pixel. Given a pair of input images ${\bf I}^{1},{\bf I}^{2}\in{\mathbb{R}}^{H\times W\times 3}$ from two different views, a weight-sharing ViT first extracts the corresponding features ${\bf F}^{1},{\bf F}^{2}$ for each view. Two parallel branches are employed to decode the geometric structures and enhance the feature alignment via cross-attention in decoder modules, following a regression head to estimate pointmaps ${\bf X}^{1,1},{\bf X}^{2,1}\in{\mathbb{R}}^{H\times W\times 3}$ along with a confidence map ${\bf C}^{1,1},{\bf C}^{2,1}\in{\mathbb{R}}^{H\times W}$ for each image view. Generally, ${\bf X}^{n,m}$ indicates the pointmap ${\bf X}^{n}$ from camera $n$ expressed in camera $m$’s coordinate frame, which is obtained by a rigid transformation:
@@
The task for Decoder 1 and its regression head estimate the 3D points for ${\bf I}^{1}$ in its own coordinate system while Decoder 2 and its regression head are responsible for estimating pixel-wise 3D coordinates for ${\bf I}^{2}$ in ${\bf I}^{1}$’s coordinate system after a rigid transformation of global rotation and translation. In the following contents, we will first introduce our POMATO with pairwise input images and then extend it to the video sequence input with our temporal motion module.
@@ 3.2 Pointmap Matching with Pairwise Input
$\mathcal{L}_{\text{m}}=\left\|\frac{1}{z_{m}}{{\bf X}^{2,1}_{m}}-\frac{1}{\bar{z}_{m}}{\bar{{\bf X}}^{2,1}}_{m}\right\|,\vskip-2.84526pt$ (3)
where $\bar{{\bf X}}^{2,1}_{m}$ is the ground truth pointmap matching, which can be obtained following Eq. 2 on the 2D tracking dataset with the depth and camera information. $z_{m},\bar{z}_{m}$ are the same norm factor defined in DUSt3R. The matching confidence ${\bf C}^{2,1}_{m}$ is also learned following the confidence loss for Head1 and Head2 within valid regions:
$\mathcal{L}_{\text{mconf}}={\bf C}^{2,1}_{m}\mathcal{L}_{\text{m}}-\alpha\text{log}{\bf C}^{2,1}_{m}\vskip-2.84526pt$ (4)
@@ 3.3 Dynamic Mask Estimation
Taking advantage of the explicit pointmap matching head, our POMATO can directly perform dynamic mask estimation without introducing an assistant module such as the optical flow model, getting rid of the additional computation cost and the potential domain gap. For an image pair $\{{\bf I}^{i},{\bf I}^{j}\}$ along with the estimation of ${\bf X}^{j,i}$ from Head2 and ${\bf X}^{j,i}_{m}$ from Head3, the dynamic mask ${\bf D}^{j,i}$ can be obtained by comparing the difference between ${\bf X}^{j,i}$ and ${\bf X}^{j,i}_{m}$:
[…]
Figure 4: Architecture of our temporal motion module. We insert a transformer-based motion module (in shallow yellow) into the vanilla DPT [33] head to enhance the temporal consistency.
3.4 Temporal Motion Module
@@
With the fundamental capability of geometric estimation and pointmap matching for pairwise images, we follow [6] and extend our POMATO to 4D video sequences by inserting a transformer-based motion module into the vanilla DPT head to construct the ”temporal DPT head”, which is illustrated in Fig.4. For a set of decoder tokens ${\bf G}\in{\mathbb{R}}^{B,T,N,C}$ where $B,T,N,C$ represent the batch size, window length of a video sequence, token number, and token dimension, respectively, we merge the token number dimension into the batch axis and apply the motion module which consists of two blocks of standard multi-head self-attention modules and feed-forward networks along the temporal dimension $T$. To reduce the computation cost, the temporal motion modules are applied to features of low resolution.
3.5 Downstream Temporal Tasks
Given a video sequence of $T$ frames ${\bf I}^{t_{1}},{\bf I}^{t_{2}},\ldots,{\bf I}^{t_{T}}$}, we construct a unique set of stereo image pairs for each task. As illustrated in Fig. 5, the flexible construction of input pairs—combined with the proposed temporal motion module and pointmap matching head—enables POMATO to seamlessly address downstream temporal tasks, including 3D point tracking, video depth estimation, and 3D reconstruction. The keyframe selection strategy and input formulation for each task are detailed in the following section.
Besides the default regression losses for Head1 and Head2, and predefined losses Eq. 3 and Eq. 4 for Head3, we further employ a temporal consistency loss, $\mathcal{L}_{\text{t}}$, which will be described in detail below.
In addition to the default regression losses for Head1 and Head2, and the predefined losses in Eq. 3 and Eq. 4 for Head3, we further introduce a temporal consistency loss, $\mathcal{L}_{\text{t}}$, which will also be described in detail below.
Figure 5: Inference pipelines for point tracking, video depth, and multi-view reconstruction. $t_{k}$ indicates the keyframe. With the help of the motion module and flexible input construction, POMATO can be easily applied to downstream temporal tasks.
3D Point Tracking. As illustrated at the top of Fig.5, the keyframe is set to the first image of the global video sequence and fed to the proposed Head3 to obtain the pointmap matching result of each query point (initialized at the first image) under the coordinate system of each reference frame $\{{\bf X}^{t_{1},t_{1}}_{m},{\bf X}^{t_{1},t_{2}}_{m},{\bf X}^{t_{1},t_{3}}_{m},...{\bf X}^{t_{1},t_{T}}_{m}\}$, while the set of reference frames $\{{\bf I}^{t_{1}},{\bf I}^{t_{2}},{\bf I}^{t_{3}},...{\bf I}^{t_{T}}\}$ is fed to the Head1 to obtain the pointmap under each ego coordinate system. The dense tracking results can be further sparsified by indexing the 2D coordinates. When inference on a video longer than $T$ frames, a simple sliding-window approach with an overlap of four frames is adopted to enhance the consistency between adjacent video windows. The temporal consistency loss $\mathcal{L}_{\text{t}}$ for tracking is:
$\mathcal{L}_{\text{t}}=\frac{1}{T}\sum_{i=1}^{T}\left\|\frac{{{\bf X}^{t_{1},t_{i}}_{m}}}{z_{m}^{T}}-\frac{{\bar{{\bf X}}^{t_{1},t_{i}}}_{m}}{\bar{z}_{m}^{T}}\right\|+\left\|\frac{{{\bf X}^{t_{i},t_{i}}}}{z^{T}}-\frac{{\bar{{\bf X}}^{t_{i},t_{i}}}}{\bar{z}^{T}}\right\|,$ (6)
where $z_{m}^{T}=$ norm $({\bf X}^{t_{1},t_{1}}_{m},{\bf X}^{t_{1},t_{2}}_{m},...,{\bf X}^{t_{1},t_{T}}_{m})$ and $\bar{z}_{T}=$ norm $(\bar{{\bf X}}^{t_{1},t_{1}}_{m},\bar{{\bf X}}^{t_{1},t_{2}}_{m},...,\bar{{\bf X}}^{t_{1},t_{T}}_{m})$. $z_{m}^{T}$ and $\bar{z}_{T}$ are defined analogously.
[…]
$\mathcal{L}_{\text{t}}=\frac{1}{T}\sum_{i=1}^{T}\left\|\frac{{{\bf X}^{t_{i},t_{i}}_{1}}}{z_{1}^{T}}-\frac{{\bar{{\bf X}}^{t_{i},t_{i}}}}{\bar{z}^{T}}\right\|+\left\|\frac{{{\bf X}_{2}^{t_{i},t_{i}}}}{z_{2}^{T}}-\frac{{\bar{{\bf X}}^{t_{i},t_{i}}}}{\bar{z}^{T}}\right\|,$ (7)
where ${\bf X}^{t_{i},t_{i}}_{1}$} and ${\bf X}^{t_{i},t_{i}}_{2}$} indicate the output from Head1 and Head2, respectively. $\bar{{\bf X}}^{t_{i},t_{i}}$ is the pointmap groundtruth.